OpenStudy (anonymous):

Given log26+log29=log2X What is the value of X?

OpenStudy (johnweldon1993):

Hint* \[\large log_c({a}) + log_c({b}) = log_c({a\times b})\]

OpenStudy (johnweldon1993):

You can use that to compress the left hand side, which will then make it very easy to see what you need to do next

OpenStudy (anonymous):

Need the answer and i will fan you.

OpenStudy (johnweldon1993):

Sorry man no answers... \[\large log(26) + log(29) = log(2x)\] \[\large log(a) + log(b) = log(a\times b)\] here: a = 26 and b = 29 Use this to simplify the left, and then you will be able to solve for 'x'

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i bet it is \[\log_2(6)+\log_2(9)=\log_2(x)\] am i right?

OpenStudy (johnweldon1993):

^Hadn't considered that! Thanks for that @misty1212

OpenStudy (anonymous):

Thank you!!!